OpenStudy (anonymous):

Write an equation for the line that is parallel to the given line and passes through the given point. y = x + 8; (4, 10) A. y = x B. y = x + 6 C. y = x – 6 D. y = x + 6

OpenStudy (anonymous):

equation will be of the form y = x + c as it is parallel to y = x +8 and it will pass through (4,10) therefore 10 = 4 + c this gives c = 6 therefore equation of line is y = x + 6 option (D)

OpenStudy (anonymous):

thank u